How would you prepare 200 mL of 20 mM NaH2PO4 buffer, pH = 7.0?
Answer #1

Molarity= Mass in grams/ molecular weight * 1000/vol req

20/1000 = x/119.98 * 1000/ 200

x= 20/1000 * 119.98 *200/1000

=0.47g of NaH2PO4 in 200mL of d.H2O

Check the pH and adjust it using either acid (HCl) or base (NaOH)